OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution. In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution. @Mathematics

OpenStudy (anonymous):

Carl = 84 Connie = 137

OpenStudy (anonymous):

Let Carl had x votes so connie had x+53 votes total number is 221 therefore x+x+53 = 221 2x = 221-53 x = 168/2 = 84